OpenStudy (anonymous):

The area of a rectangular athletic field is represented by the expression48x^4+32x^3-72x square meters. Write an algebraic expression to represent one possible set of dimensions (in the sense “length times width”) of the athletic field. Include correct units with your solution.

myininaya (myininaya):

A=48x^4+32x^3-72x A=x(48x^3+32x^2-72) so length could=x width could=48x^3+32x^2-72

myininaya (myininaya):

but there are other possibilities

OpenStudy (anonymous):

Should we factor this out?

OpenStudy (anonymous):

8x(-9 + 4x2 + 6x3

OpenStudy (anonymous):

8x(6x^3 + 4x^2 -9)

OpenStudy (anonymous):

8x(-9 + 4x^2 + 6x^3)

OpenStudy (anonymous):

which South do you attend? online or on campus?